OpenStudy (18jonea):

Determine the number and type of complex solutions and possible real solutions for each of the following equations. 1) 2x^2 + 5x + 3 = 0 2) 4x^3 – 12x + 9 = 0 3) 2x^4 + x^2 – x + 6 = 0

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (18jonea):

@mathmale

OpenStudy (18jonea):

@robtobey

OpenStudy (18jonea):

I have no idea where to even start

OpenStudy (18jonea):

@Hero

OpenStudy (mathmale):

"1) 2x^2 + 5x + 3 = 0 " is a quadratic equation in the form ax^2 + bx + c=0. By comparing these two equations, determine the values of a, b and c. a= b= c = Now calculate the value of the "discriminant." This is defined as follows:\[discriminant=b^2-4ac\]

OpenStudy (18jonea):

a =2 b=5 c=3

OpenStudy (18jonea):

25-24=1

OpenStudy (18jonea):

@mathmale

OpenStudy (18jonea):

@mathmale

OpenStudy (mathmale):

Sorry for the delay in my response. You did well in calculating the discriminant. Here are the rules: 1. If the discriminant is positive, you have two real, different roots (e. g., x=4 and x=5) 2. If the discriminant is zero, you have two real, equal roots (e. g., x=2, x=2) 3. If the disc. is negative, you have two complex roots (e. g., x=2+i, x=2-i) Which one of these cases do you have in this problem? How would you describe the roots of your quadratic equation?

OpenStudy (18jonea):

there are two real roots

OpenStudy (18jonea):

now how would i find the roots? @mathmale

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (michele_laino):

for the first equation, it is suffice to apply the sandard formula: \[\Large x = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]

OpenStudy (mathmale):

Yes. Note, @18jonea, that you have already found b^2 - 4ac: it is 1. So, subst. 5 for b in Michele's equation, above, and 2 for a. then evaluate the two roots. Please share your work here.

OpenStudy (mathmale):

\[x=\frac{ -b \pm \sqrt{1} }{ 2a }\]

OpenStudy (18jonea):

\[x=-5\sqrt{1}\div4\]

OpenStudy (mathmale):

Of course, simplify Sqrt(1) to just plain 1.

OpenStudy (18jonea):

-5+1=-4/4=-1

OpenStudy (mathmale):

Here, b=4 and a=2. We must write the formula for x using parentheses: x= [-5 plus or minus 1] / 4

OpenStudy (18jonea):

-5-1=-6/4= -3/2

OpenStudy (mathmale):

Try it, please. What is -5 plus 1? What is -5 minus 1? do this work BEFORE attempting to divide by 4.

OpenStudy (mathmale):

Also, after writing one equation, move DOWN (not to the right) to type your next result. Wrong: -5-1=-6/4= -3/2 Right: (-5-1)/4 = -6/4 = -3/2 (answer)

OpenStudy (18jonea):

The roots are -1 and -3/2

OpenStudy (18jonea):

is that correct?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (mathmale):

Nice work, 18-J!

OpenStudy (mathmale):

thank you, Michele!

OpenStudy (michele_laino):

:)

OpenStudy (18jonea):

Ok Thank you so much could y'all help me with the other two as well please?

OpenStudy (mathmale):

I know you have 2 more problems. Please start a new post for each, OK? Thanks.

OpenStudy (18jonea):

Ok